# Centro de Investigación, Desarrollo e Innovación en Ingeniería Mecatrónica Universidad Católica Boliviana “San Pablo”, La Paz, Bolivia

### v1

@@ Abstract
Index Terms:
Wildfire detection, real time monitoring, deep reinforcement learning, deep learning, edge computing
@@ I Introduction
In recent years, a surge in forest fires has plagued the globe, fueled by a complex interplay of factors including climate change, deforestation, land use changes, human activities, and prolonged droughts. This escalating crisis has had a particularly devastating impact on Bolivia, where wildfires in 2024 alone have scorched over 10 million hectares (24.7 million acres) of land [1]. This staggering figure surpasses previous years, marking a significant environmental crisis for the country [2]. While developed nations utilize sophisticated early wildfire warning systems like ALERTCalifornia [3], NASA’s FIRMS program [4], and the European Forest Fire Information System [5], their replication in developing countries remains a challenge due to financial constraints.
The longer a forest fire lasts, the more difficult and costly it becomes to extinguish. Therefore, early detection of forest fires is paramount in minimizing the environmental and socioeconomic damage they inflict, which in turn requires a low rate of false alarms. Fortunately, recent advances in embedded systems, artificial intelligence (AI), and the Internet of Things (IoT) are significantly enhancing the automation of wildfire monitoring and management. Novel real-time fire detection systems utilize a combination of sensors, microphones, and cameras to identify early fire indicators like smoke, heat, sound, and gas emissions [6, 7, 8, 9].
@@ II Background
Figure 1: Overview of the ForestProtector architecture, designed for efficient wildfire monitoring.
Conventional fire detection methods, often relying on human observation from watchtowers or using old tools such as the Osborne fire finder [10], are inherently inefficient and prone to human error. Over time, different wildfire detection technologies have emerged, with advantages and disadvantages involved. These solutions can be clustered into three main categories: remote sensing-based, optical sensor-based, and IoT sensor-based. Remote sensing (e.g. from Low Earth Orbit (LEO) satellites) is particularly valuable for monitoring large forest areas at a continental or global scale [11]. Although higher resolution and frequency in satellite imagery generally translate to increased costs, ongoing advancements in satellite technology and sensor development are improving revisit times and accessibility for wild fire detection [12]. This trend points towards greater accessibility and wider use of satellite-based fire detection in developing countries in the years to come.
Unlike the optical sensors used in remote sensing, ground-based optical sensors are widely available and have been adopted for a wide range of applications due to their affordability and ease of use. These sensors can be deployed in diverse locations, including ground level, watchtowers, poles, and even unmanned aerial vehicles [13, 14]. Optical sensors fall into two main categories, RGB cameras and infrared (IR) cameras, with IR encompassing short-wave, middle-wave, and long-wave [15]. All of them ease the detection of fire traits, such as smoke or flames during day and night time. Furthermore, 360° cameras (also known as PTZ cameras) are frequently used [16], despite their higher current cost.
Wireless sensor networks (WSNs), which can be integrated into the Internet of Things (IoT), can collect data on various factors, such as temperature, humidity, pressure, and gas levels (e.g., carbon monoxide and carbon dioxide). Indeed, some investigations have used this approach to continuously calculate fire risk [6, 7]. However, as with optical-based sensing, WSNs face limitations in covering large areas due to the need for multiple nodes to achieve comprehensive monitoring [17]. On the other hand, WSNs can be combined with other sensing modalities for prompt fire detection. For example, Peruzzi et al. [8] embedded convolutional neural networks (CNNs) on low-power IoT devices for audio and image data classification in forest areas. The combined use of the CNNs provided higher accuracy (96.15%) for normal or fire classification.
@@ V Wildfire Detection Using Computer Vision
The proposed computer vision system utilizes a Raspberry Pi Camera Module v2.1 installed in the gateway, enabling the collection of RGB images with a resolution of 1920px x 1080px at 30 frames per second. During the day, the system prioritizes smoke detection over fire detection, as smoke is a more visible wildfire indicator from long distances. Inspired by our successful results classifying video on embedded devices (Jetson Nano development cards) [18], we implemented 3DCNN models for smoke recognition. These models were trained using a combination of real and synthetic videos of smoke plumes and wildfire captured at various distances.
For nighttime fire detection, the system relies on a modified version of the algorithm proposed by Gunay et al. in [19]. Our implementation prioritizes bright region detection (BRD), average magnitude difference function (AMDF), background subtraction, and color thresholding. We collected 50 night videos with fire and 50 night videos without fire. Through experimentation, we identified that 180 and 0.2 were suitable threshold values for the BRD and AMDF components, respectively. In the following sections, we describe the development of the 3DCNN-based model for daytime wildfire detection.
@@ V-A Dataset preparation
To collect the dataset detailed in Table II, we used video editing tools, scraped videos from social media websites, and used generative AI tools. First, we used After Effects to generate smoke effects with varying wind and orientation, which were then fused with videos of forests recorded in La Paz, Bolivia. We also scraped videos from social media using search terms like “wildfire,” “burning forest,” and “wild forest.” Finally, we used the AI tool Haiper [20] to generate wildfire videos from static images.
@@ VI-A Deep Reinforcement Learning Model
We trained the model using the following hyperparameters: learning rate $\alpha=0.001$, discount factor $\gamma=0.75$ (to balance the importance of immediate and future rewards), and rewards $r=5$ for correctly identifying the IoT node with the strongest signal, and $r=-1$ otherwise. To enhance stability during training, an experience replay buffer of size 100,000 was employed to store past transitions. At each step, a mini-batch of 64 transitions was randomly sampled. The epsilon-greedy policy enabled a balance between exploration and exploitation, with $\epsilon$ initialized at 1.0 and decaying at a rate of 0.995 per episode, down to a minimum of 0.01. As illustrated in Figure 4, the model demonstrated significant improvement in its decision-making capability over the episodes. Additionally, the following weights were assigned to the sensor data: $W_{s}=0.6$ (highest priority, as it provides a continuous measure strongly correlated with wildfire risk), $W_{t}=0.3$ (temperature sensor contributes significantly when combined with humidity), and $W_{h}=0.05$ (humidity sensor has the lowest individual weight).
@@ VI-B Smoke Detection
III provides a comparative analysis of the five 3DCNN architectures (S1 to S5) experimented in the project. Our experiments demonstrated the importance of incorporating modules like Average Pooling, bidirectional LSTM layers, and grayscale conversion for enhanced performance. S5 achieved the highest metrics (accuracy = 0.925, F-score = 0.93), demonstrating superior performance. Figure 5 shows the accuracy and loss over epochs, highlighting S5’s convergence and stability. Furthermore, Figure 6 shows classification examples obtained using S5, where it is worth noting that the model misclassified fog as smoke in the second sample. In the inference stage, the video captured by the camera is divided into an 8x4 grid, and each cell is analyzed individually.

### v2

# Sergio Fernandez-Testa , and Misael Mamani

@@ I Introduction
In recent years, a surge in forest fires has plagued the globe, fueled by a complex interplay of factors including climate change, deforestation, land use changes, human activities, and prolonged droughts. This escalating crisis has had a particularly devastating impact on Bolivia, where wildfires in 2024 alone have scorched over 10 million hectares (24.7 million acres) of land [deutschewelle2024]. This staggering figure surpasses previous years, marking a significant environmental crisis for the country [copernicus2024]. While developed nations utilize sophisticated early wildfire warning systems like ALERTCalifornia [alertcalifornia2024], NASA’s FIRMS program [nasa2024], and the European Forest Fire Information System [effis2024], their replication in developing countries remains a challenge due to financial constraints.
The longer a forest fire lasts, the more difficult and costly it becomes to extinguish. Therefore, early detection of forest fires is paramount in minimizing the environmental and socioeconomic damage they inflict, which in turn requires a low rate of false alarms. Fortunately, recent advances in embedded systems, artificial intelligence (AI), and the Internet of Things (IoT) are significantly enhancing the automation of wildfire monitoring and management. Novel real-time fire detection systems utilize a combination of sensors, microphones, and cameras to identify early fire indicators like smoke, heat, sound, and gas emissions [dampage2022, benzekri2020, peruzzi2023, li2024].
@@ II Background
Figure 1: Overview of the ForestProtector architecture for efficient wildfire monitoring.
Conventional fire detection methods, often relying on human observation from watchtowers or using old tools such as the Osborne fire finder [ororatech2024], are inherently inefficient and prone to human error. Over time, different wildfire detection technologies have emerged, with advantages and disadvantages involved. These solutions can be clustered into three main categories: remote sensing-based, optical sensor-based, and IoT sensor-based. Remote sensing (e.g. from Low Earth Orbit (LEO) satellites) is particularly valuable for monitoring large forest areas at a continental or global scale [chen2024]. Although higher resolution and frequency in satellite imagery generally translate to increased costs, ongoing advancements in satellite technology and sensor development are improving revisit times and accessibility for wild fire detection [yang2024]. This trend points towards greater accessibility and wider use of satellite-based fire detection in developing countries in the years to come.
Unlike the optical sensors used in remote sensing, ground-based optical sensors are widely available and have been adopted for a wide range of applications due to their affordability and ease of use. These sensors can be deployed in diverse locations, including ground level, watchtowers, poles, and even unmanned aerial vehicles [mukhiddinov2022, saydirasulovich2023]. Optical sensors fall into two main categories, RGB cameras and infrared (IR) cameras, with IR encompassing short-wave, middle-wave, and long-wave [ciprian2023]. All of them ease the detection of fire traits, such as smoke or flames during day and night time. Furthermore, 360° cameras (also known as PTZ cameras) are frequently used [shah2021], despite their higher current cost.
Wireless sensor networks (WSNs), which can be integrated into the Internet of Things (IoT), can collect data on various factors, such as temperature, humidity, pressure, and gas levels (e.g., carbon monoxide and carbon dioxide). Indeed, some investigations have used this approach to continuously calculate fire risk [dampage2022, benzekri2020]. However, as with optical-based sensing, WSNs face limitations in covering large areas due to the need for multiple nodes to achieve comprehensive monitoring [lloret2009]. On the other hand, WSNs can be combined with other sensing modalities for prompt fire detection. For example, Peruzzi et al. [peruzzi2023] embedded convolutional neural networks (CNNs) on low-power IoT devices for audio and image data classification in forest areas. The combined use of the CNNs provided higher accuracy (96.15%) for normal or fire classification.
@@ V Wildfire Detection Using Computer Vision
The proposed computer vision system utilizes a Raspberry Pi Camera Module v2.1 installed in the gateway, enabling the collection of RGB images with a resolution of 1920px x 1080px at 30 frames per second. During the day, the system prioritizes smoke detection over fire detection, as smoke is a more visible wildfire indicator from long distances. Inspired by our successful results classifying video on embedded devices (Jetson Nano development cards) [fernandez2024], we implemented 3DCNN models for smoke recognition. These models were trained using a combination of real and synthetic videos of smoke plumes and wildfire captured at various distances.
For nighttime fire detection, the system relies on a modified version of the algorithm proposed by Gunay et al. in [gunay2009]. Our implementation prioritizes bright region detection (BRD), average magnitude difference function (AMDF), background subtraction, and color thresholding. We collected 50 night videos with fire and 50 night videos without fire. Through experimentation, we identified that 180 and 0.2 were suitable threshold values for the BRD and AMDF components, respectively. In the following sections, we describe the development of the 3DCNN-based model for daytime wildfire detection.
@@ V-A Dataset preparation
To collect the dataset detailed in Table II, we used video editing tools, scraped videos from social media websites, and used generative AI tools. First, we used After Effects to generate smoke effects with varying wind and orientation, which were then fused with videos of forests recorded in La Paz, Bolivia. We also scraped videos from social media using search terms like “wildfire,” “burning forest,” and “wild forest.” Finally, we used the AI tool Haiper [haiperai2024] to generate wildfire videos from static images.
@@ VI-A Deep Reinforcement Learning Model
We trained the model using the following hyperparameters: learning rate $\alpha=0.001$, discount factor $\gamma=0.75$ (to balance the importance of immediate and future rewards), and rewards $r=5$ for correctly identifying the IoT node with the strongest signal, and r=1r=-1 otherwise. To enhance stability during training, an experience replay buffer of size 100,000 was employed to store past transitions. At each step, a mini-batch of 64 transitions was randomly sampled. The epsilon-greedy policy enabled a balance between exploration and exploitation, with $\epsilon$ initialized at 1.0 and decaying at a rate of 0.995 per episode, down to a minimum of 0.01. As illustrated in Figure 4, the model demonstrated significant improvement in its decision-making capability over the episodes. Additionally, the following weights were assigned to the sensor data: $W_{s}=0.6$ (highest priority, as it provides a continuous measure strongly correlated with wildfire risk), $W_{t}=0.3$ (temperature sensor contributes significantly when combined with humidity), and $W_{h}=0.05$ (humidity sensor has the lowest individual weight).
@@ VI-B Smoke Detection
Table III provides a comparative analysis of the five 3DCNN architectures (S1 to S5) experimented in the project. Our experiments demonstrated the importance of incorporating modules like Average Pooling, bidirectional LSTM layers, and grayscale conversion for enhanced performance. S5 achieved the highest metrics (accuracy = 0.925, F-score = 0.93), demonstrating superior performance. Figure 5 shows the accuracy and loss over epochs, highlighting S5’s convergence and stability. Furthermore, Figure 6 shows classification examples obtained using S5, where it is worth noting that the model misclassified fog as smoke in the second sample. In the inference stage, the video captured by the camera is divided into an 8x4 grid, and each cell is analyzed individually.